One number is 6 less than a second number. Twice the second number is 40 more than 4 times the first. Find the two numbers.
Answer by rapaljer(4667)

Let x = second number (since this is at the end of the first sentence!)
x-6 = the other number (the first number)

Now, substitute into the sentence:
Twice the second number IS 40 more than 4 times the first number.
2( ) = 4*( ) + 40
2(x) = 4*(x-6) + 40
2x = 4x - 24 + 40
2x = 4x + 16

Subtract 4x from each side:
2x-4x =4x-4x+16
-2x=16

Divide both sides by -2:
x=-8 = second number
x-6= first number
-8-6=-14 = first number

Check: "Twice the second number is 40 more than 4 times the first."
2*-8=4*(-14)+40
-16=-56+40
-16=-16 It checks!!

For additional help with ALL types of WORD PROBLEMS, please see my own website. Click on my tutor name "rapaljer" anywhere in algebra.com. Look for the link to my website there. From my Homepage, look for the link "Basic, Intermediate, and College Algebra: One Step at a Time", click on "Basic Algebra," and then "Chapter 1." Word Problems can be found in "Sections 1.09 and 1.10." See also the solutions of many of these exercises in the "MATH IN LIVING COLOR" pages, where the problems are solved in COLOR. My own students and others have been unanimous in telling me that my own explanations are easier to understand than the expensive textbooks that are written for mathematicians and publishers!! MY EXPLANATIONS WERE WRITTEN FOR STUDENTS WHO HAVE A HARD TIME UNDERSTANDING MATH!! And they are all FREE!!

I hope you like my website!!! If you do, please recommend it to your family and friends!! At least check out the JOKE PAGE--all safe jokes for the little ears in your family!!

Dr. Robert J. Rapalje, Retired
Seminole State College of Florida
Altamonte Springs Campus